# Aligning Human and LLM Judgments: Insights from EvalAssist on Task-Specific Evaluations and AI-assisted Assessment Strategy Preferences

### v1

@@ 6.2. Alignment: Refining Criteria, Changing the AI Evaluator, and Altering Expected Results
We observed various user interactions throughout the evaluation process that mirror principles seen in interactive machine learning (iML), where users iteratively refine models by selecting examples, labeling data, and evaluating outputs (Patel et al., 2010; Simard et al., 2014). In line with iML approaches, users in our study adjusted their criteria throughout the evaluation, making them more general or specific as needed. This adaptive behavior reflects the need for flexible interfaces that allow seamless transitions between specifying, revising, and evaluating criteria. As noted in iML, reducing user effort while improving output alignment is crucial (Desmond et al., 2022). In some cases, midway through the evaluation, users even changed their expected outcomes to align more closely with the AI’s judgments—especially after reading the explanations, which often convinced them of the AI’s output. Users frequently revised their criteria, whether by adding or removing scales in the rubric condition, or rephrasing their criteria in the pairwise condition. They also experimented with different AI evaluators, switching between models like Mixtral and LLaMA to compare results. This suggests that the path to ”alignment” can take many forms, from adjusting one’s own expectations, to switching the AI model making the judgments, to modifying the criteria itself in search of a better fit. This process could be better supported in the future by automatically testing criteria with multiple models and showing results side-by-side.
@@ 6.3. Adaptive Evaluation Strategies: Balancing Clarity and Flexibility
Participants expressed a preference for Direct Assessment due to the clarity and control it offered, allowing for detailed, individualized evaluations. On the other hand, Pairwise Assessment was valued for its flexibility, particularly in more subjective tasks. Some participants preferred a hybrid approach, selecting an evaluation method based on the specific task. The higher number of evaluations in direct assessment suggests that users felt more engaged, likely because it provided them with a greater sense of control. This aligns with the reasons participants cited for favoring direct assessment. These findings highlight the importance of adaptable evaluation strategies tailored to task type. AI systems could benefit from offering multiple evaluation modes or an adaptive assessment strategy that adjusts based on task complexity. For example, users might opt for direct assessment for objective tasks and switch to pairwise ranking for more subjective ones. Additionally, some users might prefer a combination approach—starting with direct assessment to filter out responses and then using pairwise ranking to further refine the best options. Future research can explore hybrid systems that adapt dynamically based on user workflows.
@@ 6.4. Bias Awareness and Explanation Visibility in Evaluation Strategies
Positional bias was considered helpful when present, prompting participants to revise and improve their criteria. This result suggests that highlighting AI biases can enhance the evaluation process and improve alignment between human and AI judgments. While our study focused on positional bias, there is an opportunity to explore the representation of other biases, such as self-enhancement bias and verbosity bias. For example, is the model ranking outputs highly because they were generated by the AI evaluator, or is it favoring longer responses even if they are less accurate or clear? Research is lacking on how users perceive such bias flagging and how it impacts their criteria refinement.
@@ 7. Conclusion
We introduced EvalAssist a tool designed to help practitioners refine evaluation criteria using both direct and pairwise assessment strategies. The tool provides positional bias metrics and explanations for each AI judgment. In a controlled experiment with 15 machine learning practitioners, we examined how users refine their criteria and identified key differences between the two evaluation approaches. Direct Assessment was preferred for its clarity and control, while Pairwise Assessment was valued for flexibility in subjective tasks. Users often refined their criteria by increasing specificity or simplifying scales to improve alignment with task needs. Human-AI alignment was stronger in the rubric-based (Direct Assessment) condition, highlighting the importance of clear evaluation criteria. Our findings suggest that AI-assisted evaluation systems should offer flexible, adaptive strategies, allowing users to switch between direct and pairwise methods. Such systems should support ongoing criteria refinement, provide clear explanations, and improve alignment between human and AI judgments to enhance evaluation reliability.

### v2

@@ 8.2. Alignment: Refining Criteria, Changing the AI Evaluator, and Altering Expected Results
We observed various user interactions throughout the evaluation process that mirror principles seen in interactive machine learning (iML), where users iteratively refine models by selecting examples, labeling data, and evaluating outputs (Patel et al., 2010; Simard et al., 2014). In line with iML approaches, users in our study adjusted their criteria throughout the evaluation, making them more general or specific as needed. This adaptive behavior reflects the need for flexible interfaces that allow seamless transitions between specifying, revising, and evaluating criteria. As noted in iML, reducing user effort while improving output alignment is crucial (Desmond et al., 2022). In some cases, midway through the evaluation, users even changed their expected outcomes to align more closely with the AI’s judgments, especially after reading the explanations, which often convinced them of the AI’s output. Users frequently revised their criteria, whether by adding or removing scales in the rubric condition, or rephrasing their criteria in the pairwise condition. They also experimented with different AI evaluators, switching between models like Mixtral and LLaMA to compare results. This suggests that the path to ”alignment” can take many forms, from adjusting one’s own expectations, to switching the AI model making the judgments, to modifying the criteria itself in search of a better fit. This process could be better supported in the future by automatically testing criteria with multiple models and showing results side-by-side. Future work can investigate the impact of different types of refinement on alignment and, if significant patterns emerge, explore tooling that nudges users toward specific refinements to improve alignment and evaluation outcomes.
@@ 8.3. Adaptive Evaluation Strategies: Balancing Clarity and Flexibility
Participants expressed a preference for direct assessment due to the clarity and control it offered, allowing for detailed, individualized evaluations. On the other hand, pairwise comparison was valued for its flexibility, particularly in more subjective tasks. Some participants preferred a hybrid approach, selecting an evaluation method based on the specific task. The higher number of evaluations in direct assessment suggests that users felt more engaged, likely because it provided them with a greater sense of control. This aligns with the reasons participants cited for favoring direct assessment. These findings highlight the importance of adaptable evaluation strategies tailored to task type. AI systems could benefit from offering multiple evaluation modes or an adaptive judgment strategy that adjusts based on task complexity. For example, users might opt for direct assessment for objective tasks and switch to pairwise ranking for more subjective ones. Additionally, some users might prefer a combination approach, starting with direct assessment to filter out responses and then using pairwise ranking to further refine the best options. Future research can explore hybrid systems that adapt dynamically based on user workflows.
8.4. Trust Across Judgment Strategies
We observed no significant differences in trust across task types or between the two AI judge strategies (direct assessment and pairwise comparison). This indicates that participants did not inherently favor one strategy over the other in terms of trustworthiness, and the design of these strategies did not introduce unintended trust disparities. While prior literature highlights how task framing and algorithmic presentation can disproportionately influence trust (Greiner et al., 2024), our findings suggest that neither judge strategy nor task type significantly impacted trust. Open-ended feedback revealed another area for exploration: the influence of bias indicators and explanations on trust. Participants noted that both bias and the quality of explanations played a role in shaping their trust in the AI evaluator. Explanations were particularly valued in the rubric condition, suggesting that future research could investigate how different explanation types affect trust across contexts. Overall, it is encouraging that no major trust differences were observed between pairwise and direct assessment strategies, indicating both are perceived as equally trustworthy by users.
@@ 8.5. Bias Awareness and Explanation Visibility in Evaluation Strategies
Positional bias was considered helpful when present, prompting participants to revise and improve their criteria. This result suggests that highlighting AI biases can enhance the evaluation process and improve alignment between human and AI judgments. While our study focused on positional bias, there is an opportunity to explore the representation of other biases, such as self-enhancement bias and verbosity bias. For example, is the model ranking outputs highly because they were also generated by the AI evaluator LLM, or is it favoring longer responses even if they are less accurate or clear? Incorporating a suite of bias indicators tailored to different contexts and tasks could strengthen users’ ability to evaluate outputs effectively. This approach aligns with the goal of fostering transparency and user trust in AI systems by equipping users with the tools to identify and mitigate potential biases.
@@ 9. Conclusion
We introduced EvalAssist a tool designed to help practitioners refine evaluation criteria using both direct and pairwise judgment strategies. The tool provides positional bias metrics and explanations for each AI judgment. In a controlled experiment with 15 machine learning practitioners, we examined how users refine their criteria and identified key differences between the two evaluation approaches. Direct assessment was preferred for its clarity and control, while pairwise comparison was valued for flexibility in subjective tasks. Users often refined their criteria by increasing specificity or simplifying scales to improve alignment with task needs. Human-AI alignment was stronger in the rubric-based (direct assessment) condition, highlighting the importance of clear evaluation criteria. While our participants found EvalAssist a valuable tool to be integrated into their work practices, findings from our experiment highlight areas for future improvements: (1) AI-assisted evaluation systems should offer flexible, adaptive strategies, allowing users to switch between direct and pairwise methods, (2) they should support ongoing criteria refinement, provide clear explanations, and improve alignment between human and AI judgments to enhance evaluation reliability, and (3) they should support involving multiple stakeholders in the criteria refinement process to pave the way to alignment.
Appendix A Additional Figures
Appendix B Direct Assessment Prompts for EvalAssist AI Evaluators
B.1. Mixtral-8x7b-instruct-v01
B.1.1. Assessment Prompt
<s> [INST] You are presented with a response generated subject to a context. The context includes information relevant to the nature or generation of the response. You will assess the quality of the response subject to an evaluation criteria.
###Context: context_variables
###Response: response
###Evaluation criteria: criteria options
Briefly assess the quality of the response subject to the evaluation criteria. Focus on the evaluation criteria during assessment, do not provide a general assessment. [/INST]
Assessment:
B.1.2. Summarization Prompt
<s> [INST] Summarize the following assessment while maintaining the pertinent details. Assessment: assessment [/INST]
Assessment Summary:
B.1.3. Answer Selection Prompt
</s> [INST] Now consider the evaluation criteria and choose a final answer.
Validate the answer against the assessment.
###Evaluation criteria: criteria options [/INST] Answer:
B.2. llama-3-8b-instruct and llama-3-70b-instruct
B.2.1. Assessment Prompt
<|begin_of_text|> <|start_header_id|>system<|end_header_id|> You are an fair and objective evaluator.<|eot_id|>
<|start_header_id|>user<|end_header_id|>
You are presented with a response generated subject to a context. The context includes information relevant to the nature or generation of the response. You will assess the quality of the response subject to an evaluation criteria.
###Context: {context_variables}
###Response: {response}
###Evaluation criteria: criteria options
Briefly assess the quality of the response subject to the evaluation criteria. Focus on the evaluation criteria during assessment, do not provide a general assessment.
Assessment: <|eot_id|>
<|start_header_id|>assistant<|end_header_id|>
B.2.2. Summarization Prompt
<|begin_of_text|>
<|start_header_id|>system<|end_header_id|>
You are a helpful assistant<|eot_id|>
<|start_header_id|>user<|end_header_id|>
Summarize the following assessment while maintaining the pertinent details.
Assessment: {assessment}
Assessment Summary:<|eot_id|>
<|start_header_id|>assistant<|end_header_id|>
B.2.3. Answer Selection Prompt
<|eot_id|> <|start_header_id|>user<|end_header_id|>
Now consider the evaluation criteria and choose a final answer. Validate the answer against the assessment.
###Evaluation criteria:
criteria
options
Answer:
<|eot_id|><|start_header_id|>
assistant
<|end_header_id|>
B.3. Prometheus-8x7b-v2
###Task Description:
A context that includes information relevant to the nature or generation of the response, a response to evaluate, and a score rubric representing an evaluation criteria are given.
1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general.
2. After writing a feedback, choose a score from the score rubric. Choose one of: {score_instructions}.
3. The output format should look as follows: "Feedback: (write a feedback for criteria) [RESULT] (Choose one of: {score_instructions})"
4. Please do not generate any other opening, closing, or explanations.
###Context: {context}
###Response to evaluate: {response}
###Score Rubrics: [{criteria}] {score_rubric}
###Feedback:
Appendix C Pairwise Assessment Prompts for EvalAssist AI Evaluators
C.1. Mixtral-8x7b-instruct-v01
C.1.1. Assessment Prompt
<s> [INST] You are provided a pair of responses (Response {option_a and Response {option_b}) generated subject to a context. You will choose the better quality response subject to the quality criteria.
This is the context: {context_variables}
This is the quality criteria:
{criteria_name}
{criteria_description}
Response {option_a}: {response_a}
Response {option_b}: {response_b}
Keeping the quality criteria in mind, briefly assess which response is better. Focus on the quality criteria during assessment, do not provide a general assessment.[/INST] Assessment:
C.1.2. Summarization Prompt
<s> [INST] Summarize the following assessment while maintaining the pertinent details. Assessment: {assessment}[/INST] Assessment Summary:
C.1.3. Answer Selection Prompt
</s> [INST] Now considering the quality criteria, which response is better quality? Validate the answer against the assessment. {options [/INST] Answer:
C.2. llama-3-8b-instruct and llama-3-70b-instruct
C.2.1. Assessment Prompt
<|begin_of_text|> <|start_header_id|>system<|end_header_id|> You are an fair and objective evaluator.<|eot_id|>
<|start_header_id|>user<|end_header_id|>
You are provided a pair of responses (Response {option_a} and Response {option_b}) generated subject to a context. You will choose the better quality response subject to the quality criteria.
This is the context: {context_variables}
This is the quality criteria:
{criteria_name}
{criteria_description}
Response {option_a}: {response_a}
Response {option_b}: {response_b}
Keeping the quality criteria in mind, briefly assess which response is better. Focus on the quality criteria during assessment, do not provide a general assessment.
<|eot_id|>
<|start_header_id|>assistant<|end_header_id|>
Assessment:
C.2.2. Summarization Prompt
<|begin_of_text|><|start_header_id|>
system<|end_header_id|>
<|eot_id|><|start_header_id|>user<|end_header_id|>
Summarize the following assessment while maintaining the pertinent details. Assessment: {assessment}
Assessment Summary:<|eot_id|>
<|start_header_id|>assistant<|end_header_id|>
Assessment Summary:
C.2.3. Answer Selection Prompt
<|eot_id|> <|start_header_id|>user<|end_header_id|>
Now consider the evaluation criteria and choose a final answer. which response is better quality? Validate the answer against the assessment.
options
<|eot_id|><|start_header_id|>assistant<|end_header_id|>
Answer:
C.3. Prometheus-8x7b-v2
###Task Description: A context that includes information relevant to the nature or generation of the response, a response to evaluate, and a score rubric representing a evaluation criteria are given.
1. Write a detailed feedback that assess the quality of two responses strictly based on the given score rubric, not evaluating in general.
2. After writing a feedback, choose a better response between Response {option_a} and Response {option_b}. You should refer to the score rubric.
3. The output format should look as follows: "Feedback: (write a feedback for criteria) [RESULT] ({option_a} or {option_b})”
4. Please do not generate any other opening, closing, and explanations.
###Context: {context_variables}
###Response {option_a}: {response_1}
###Response {option_b}: {response_1}
###Score Rubric: {rubric}
###Feedback: